• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Reimer

    Log entry time 17:30:45 on May17,2008

    Entry number 238462

    This entry is a followup to: 238459

    keyword=Target heat load OK

    The target heat load appears to be OK now that we have beam back.


    FIGURE 1